OpenStudy (anonymous):

How do you integrate this using substitution ? ∫x+2/X^2+4x dx , with u=x^2+4x

OpenStudy (lgbasallote):

if u = x^2 + 4x then du = 2x + 4 agree?

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

then du = 2(x+2) do you see it?

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

so can you see the substitution now?

OpenStudy (anonymous):

yea thanks :) I was trying to solve it without factoring.How do you know when to factor it ?

OpenStudy (lgbasallote):

experience? lol idk. i was looking at the numerator. it say x+ 2 and you said solve by substitution so i was just looking on how to make it look like x+2

OpenStudy (anonymous):

oh okay.Thanks once again

OpenStudy (lgbasallote):

<tips hat>